OpenStudy (anonymous):

2sqrt{45}-6sqrt{5} __how do i simplify this radical expression

OpenStudy (anonymous):

hint : 45 = 9*5

OpenStudy (anonymous):

first part of answer \[\sqrt{45}=\sqrt{9\times 5}=\sqrt{9}\times \sqrt{5}=3\sqrt{5}\]